### v1

# Mapping Bias in Vision Language Models: Signposts, Pitfalls, and the Road Ahead

Abstract
As Vision Language Models (VLMs) gain widespread use, their fairness remains under-explored. In this paper, we analyze demographic biases across five models and six datasets. We find that portrait datasets like UTKFace and CelebA are the best tools for bias detection, finding gaps in performance and fairness between LLaVa and CLIP models. However, scene based datasets like PATA, VLStereoSet fail to be useful benchmarks for bias due to their construction. As for pronoun based datasets like VisoGender, we receive mixed signals as only some subsets of the data are useful in providing insights. To alleviate this problem, we introduce a more difficult version of VisoGender to serve as a more rigorous evaluation. Based on these results, we call for more effective and carefully designed datasets to ensure VLMs are both fair and reliable. [^1] Code for analysis and evaluation can be found at https://github.com/KuleenS/VLMBiasEval
Mapping Bias in Vision Language Models: Signposts, Pitfalls, and the Road Ahead
Kuleen Sasse1,2, Shan Chen3,4,5, Jackson Pond3,4, Danielle S. Bitterman3,4,5*thanks: Co-Senior Authors: Danielle S. Bitterman, John Osborne, John Osborne211footnotemark: 1 1Johns Hopkins University, 2University of Alabama at Birmingham, 3Harvard Medical School, 4Mass General Brigham, 5Boston Children’s Hospital Correspondence: ksasse@uab.edu
1 Introduction
Vision Language Models (VLMs) have demonstrated human and even superhuman performance across a wide range of tasks, including image captioning, visual question answering (VQA), and object detection Chen et al. (2024); Liu et al. (2023); Radford et al. (2021). Despite their impressive capabilities, these models exhibit harmful biases that can negatively impact user experience, perpetuate inequities, and lead to potential legal consequences Kim (2022); Zhang et al. (2020). While biases and fairness in Large Language Models (LLMs) have been extensively studied, with numerous datasets created to support these efforts Gallegos et al. (2023), fairness and bias in VLMs have been less explored, with fewer available datasets and less thorough evaluation Lee et al. (2023).
In this work, we conduct a comprehensive analysis of bias in VLMs, evaluating both the models and the datasets used to assess those biases. Our contributions are threefold:
  • We analyze the biases present in five prominent VLMs across six datasets, providing new insights into model fairness.
  • We highlight the limitations of current bias evaluation datasets, demonstrating that many fail to capture the full scope of bias in multimodal models.
  • We propose future directions for developing more comprehensive and effective bias evaluation methods for VLMs, focusing on improving dataset design to mitigate conflicts between visual and textual inputs.
With this work, we aim to improve the understanding of bias in VLMs and to provide a foundation for more robust and fair models in multimodal tasks.
Figure 1: VLMs Bias Benchmark Taxonomy
2 Related Work
Bias in LLMs has been extensively studied, especially as these models have scaled in both capability and scope. However, as VLMs continue to evolve, their corresponding benchmarks have largely followed similar developmental frameworks as those used for LLMs. In this section, we review key studies on bias in LLMs and VLMs and categorize existing evaluation benchmarks based on biases related to portraits, scenes, and pronouns in VLMs.
2.1 Bias in LLMs
LLMs have made significant advancements, yet they continue to inherit and sometimes amplify societal biases, particularly in open-ended language generation tasks. Several benchmarks have been developed to quantify and mitigate these biases. For instance, the BOLD benchmark Dhamala et al. (2021) assesses social biases across categories such as gender, race, and religion in models’ free-form text generation. Similarly, Parrish et al. (2022) introduced the BBQ benchmark, which measures biases in multiple-choice question-answering tasks, focusing on how LLMs handle stereotypical or biased scenarios. StereoSet Nadeem et al. (2020) evaluates stereotypical biases across domains such as gender, profession, race, and religion by presenting models with stereotypical and non-stereotypical sentence pairs to reveal the model’s inherent biases.
Expanding on these efforts, CrowS-Pairs Nangia et al. (2020) evaluates bias in sentence pairs across nine categories, examining how models select between socially biased and neutral sentences. HONEST Nozza et al. (2021) adopts a multilingual approach, analyzing harmful sentence completions to understand how LLMs generate biased or hurtful content across different languages. While these benchmarks primarily focus on text-based biases, they offer valuable insights into how LLMs may introduce similar biases when combined with visual data in VLMs.
2.2 Bias in VLMs
VLMs, which integrate visual and textual data, present a uniquely complex landscape for bias evaluation. Unlike purely text-based models, VLMs such as CLIP and DALL-E are susceptible to biases inherent in both the images and the text used during training. This dual-source bias amplifies and complicates fairness concerns, as prejudices from both modalities can influence the model’s behavior. Several studies have explored how different pretraining methods and data mixtures impact VLM performance and bias propagation in downstream tasks Wang et al. (2022b); Fabbrizzi et al. (2023); Cabello et al. (2023). These studies consistently highlight how VLMs can reinforce societal stereotypes related to gender, race, and professional roles.
For example, Wang et al. (2022a) introduced the OccupationBias dataset to assess how VLMs perpetuate occupational stereotypes. Their findings show that VLMs often associate certain professions with specific genders or races, independent of contextual information. Similarly, Fraser and Kiritchenko (2024) introduced the PAIRS dataset, which presents parallel images of individuals from different racial and gender backgrounds to evaluate how VLMs adjust outputs based on visual cues. Their work demonstrates that models frequently modify predictions and descriptions based on superficial attributes like race or gender, even when the scene remains unchanged.
Another notable dataset in this area is VisoGender Hall et al. (2023), which examines gender bias in pronoun resolution. VisoGender tests whether models correctly resolve pronouns in images of people performing professional roles, revealing significant gender biases in multimodal models. For instance, when describing a doctor and a patient, models often default to gendered pronouns based on societal stereotypes rather than visual evidence. These findings underscore the broader challenge of mitigating bias in multimodal models, where both textual and visual information contribute to predictions.
2.3 Categorization of VLM Bias Benchmarks
As shown in Figure 1, existing benchmarks for VLMs can be categorized based on the types of biases they target: portrait-based, scene-based, and pronoun-based biases. Recent research has predominantly focused on portrait-based biases in VLMs Wu et al. (2024); Han et al. (2024). Our work contributes to these efforts by examining multiple benchmarks within an organized taxonomy and identifying limitations that may lead to inaccurate measurements of a VLM’s true bias risk.
Portrait-based biases address biases related to individual appearances, such as facial features, race, gender, and age. Datasets like CelebA Liu et al. (2015) and UTKFace Zhang et al. (2017) are commonly used to evaluate these biases. CelebA, a large-scale portrait dataset, includes attributes such as "blond hair" and "heavy makeup" and is widely used to study gender bias in facial recognition models. UTKFace includes labels for age, race, and gender and is used to evaluate how consistently models identify these attributes across demographic groups.
Scene-based biases evaluate how VLMs interpret complex environments involving multiple objects or people, and how bias influences their understanding of these scenes. Datasets like VLStereoSet Tian et al. (2022) and PATA Singh et al. (2022) challenge models with images accompanied by stereotypical and anti-stereotypical captions. VLStereoSet, for example, tests whether models align stereotypical captions with particular demographics, revealing how biases manifest in everyday visual scenes. PATA extends this by including both positive and offensive captions, investigating whether models generate biased outputs based on protected categories like race or gender.
Pronoun resolution-based biases focus on gender bias in resolving pronouns, particularly in professional or occupational contexts. VisoGender is a prominent benchmark in this category Hall et al. (2023), designed to test gender bias when resolving pronouns in images of people performing professional roles. For instance, VisoGender presents images of doctors or engineers, requiring models to select the correct pronouns without relying on stereotypical gender associations. An extension, VisoGender Adversarial, introduces misleading gender labels to test model robustness against biased assumptions, forcing the model to rely solely on visual evidence rather than on stereotypical gender roles.
3 Methods
Dataset Prompt Labels
CelebA Does the person in the photo have blond hair? Yes, No
Does the person in the photo have heavy makeup?
PATA Which one is the correct caption of this image? Multiple Choice Question
UTKFace What is the race of the person in the picture?
What is the age of the person in the picture?
What is the gender of the person in the picture?
VisoGender Which one is the correct caption of this image?
VLStereoSet Which one is the correct caption of this image?
Table 1: Prompt templates used in our experiments.
3.1 Datasets
We consider five different image datasets that include protected attributes like sex, race, religion, and age to evaluate the fairness of these models: CelebA, PATA, UTKFace, Visogender, and VLStereoSet.
3.1.1 CelebA
CelebA is a large-scale portrait dataset consisting of 200k celebrities faces with 40 different attributes apparent in the portrait, such as sex, the presence of heavy makeup or glasses in the portrait. Similarly to Kim et al. Kim et al. (2024), we focus on classifying two attributes: blond hair and heavy makeup. We use the provided test set of the model to run our evaluation on.
3.1.2 PATA
PATA is a set of about 5,000 public images crawled from the web that have been organized into 24 scenes. Each image is labeled with sex (male or female), race (White, Black, Indian, East-Asian, and Latino-Hispanic), and two age groups (youth and adults). Each image has two sets of captions associated with it: One set of positive captions with desirable descriptions of the person and one set of offensive or untoward text. These captions are developed with the protected attributes in mind and try to elicit a biased response from the model. For example, captions describing a person whose gender is the protected attribute could be a "photo of a smart engineer", trying to elicit a common bias that women cannot be smart and/or engineers. We use the entire dataset to run our evaluation as no training or test split is provided.
3.1.3 UTKFace
UTKFace is another portrait dataset that consists of over 20K portraits in the wild that cover a large variety of poses, facial expressions, illumination, etc. Each image is labeled with age, race, and sex. Similarly to Yang et al. (2024), we bin the age label into 5 categories: "Child" (020)020(0-20)( 0 - 20 ), "Young" (2040)2040(20-40)( 20 - 40 ), "Middle-Aged" (4060)4060(40-60)( 40 - 60 ), and "Senior" (60)limit-from60(60-)( 60 - ). We focus on classifying all three attributes and measure the bias based on the other two. We call each dataset using a two part name. The first part is what the model is predicting. The second part is what the bias is measured over. For example, "UTKFace Race Gender" would be predicting the race of the person while measuring the difference in performance across genders. As there is no test set provided, we sample 2,000 photos of the 20,000 as a test set to run our evaluation on.
3.1.4 VisoGender
VisoGender is a set of about 700 images that are used to benchmark gender bias in VLMs. The dataset evaluates the ability of a model to correctly resolve the pronoun of a person in a picture of them working in a professional role. The dataset measures how much the model associates different occupations with different genders. They split these images into two smaller datasets: Occupation-Object (OO) and Occupation-Participant (OP). The OO dataset has only one person working in a profession in the image, and the model must choose that person’s pronoun. The OP dataset has two people in the image: one who is working in a profession and one who is participating. The task is for the model must resolve the pronoun of the one who is working. We use the entire dataset to run our prediction as no train or test split is provided.
3.1.5 VLStereoSet
VLStereoSet is a set of about 1000 images split between 4 categories: Gender, Profession, Race, and Religion. Each image is given both a stereotypical and anti-stereotypical caption. This dataset aims to to measure the association with a certain social group with stereotypical attributes. We use the entire dataset as no train or test split is provided.
3.2 Models
We evaluate the performance and biases of various VLMs across different sizes and architectures, including both open- and closed-source models. For the open-source models, we assess the LLaVa v1.6 series, specifically: llava-v1.6-34b-hf, llava-v1.6-vicuna-7b, and llava-v1.6-mistral-7b-hf Liu et al. (2023). Additionally, we evaluate two CLIP models, CLIP L-224 and CLIP L-336, with the latter serving as the Vision Transformer (ViT) component for all the LLaVa models mentioned. For the closed-source models, we include Google’s May 2024 release: gemini-1.5-flash-001 Radford et al. (2021); Goldin et al. (2024).
All experiments were conducted on Nvidia A100 80GB GPUs, with CUDA version 12.1, transformers version 4.41.0, and the API models run through Google Vertex AI with google-cloud-platform version 1.60.0. For reproducibility, we set the temperature to 0 in all cases.
Figure 2: Diagram showing the process of creating the adversarial version of the VisoGender Dataset
3.3 Prompts
The prompts used in the experiments (shown in Table 1) were adapted from those recommended by the LLaVa authors for model evaluation and inspiration The initial questions for each prompt were inspired by prompts from VQA datasets like VQAv2 Goyal et al. (2017), GQA Hudson and Manning (2019), and MME Fu et al. (2024) datasets. We manually tested and refined these prompts to ensure that the models produced accurate answers on a small subset of training data.
For binary Yes/No questions, we ask the model the relevant question and append the recommended LLaVa prompt for binary questions: “Answer the question using a single word or phrase.”
For multiple-choice questions(MCQ), we ask the model to choose the correct caption for the image from a list of options, ending wit hrecommended prompt for MCQ type questions for LLaVa: “<MCQ Options> Answer with the option’s letter from the given choices directly.”
3.4 Bias Dataset Effectiveness Experiment
For prediction generation, we allow models to produce a distribution over the next possible tokens. From this distribution, we select the most probable token corresponding to either ("Yes", "No") for binary questions and ("A", "B", "C", etc.) for multiple-choice questions using a process similar to the lm-evaluation-harness by EleutherAI Gao et al. (2023).
To evaluate how effectively the datasets measure bias in response to visual input, we run each dataset twice: once with the image and once without the image. This approach helps us determine whether the dataset is an effective tool for measuring bias and performance in vision tasks. We hypothesize that if the dataset is indeed effective, there should be a significant performance and fairness gap between the two runs — with much better performance and fairness when the image is included.
To compare our models’ effectiveness, we include a random classifier baseline. This baseline uniformly samples an answer from the label space for each question.
3.5 Vision Reliance Experiment
To assess model robustness, we re-evaluate performance using a adversarial version of the VisoGender dataset, as shown in Figure 2. In this experiment, we modify the prompts to intentionally mislead the model by providing incorrect gender descriptions for the individuals in the image. Specifically, we append an inaccurate gender description to the end of the question. For example, if the image shows a male doctor attending to a female patient, the modified prompt might read: "Which one is the correct caption of this image of a woman helping a man?"
Despite this adversarial input, we expect the model to identify the correct caption, such as "the doctor is chatting with his patient," rather than being misled by the incorrect gender description. This experiment tests the model’s ability to rely on visual evidence rather than textual biases.
3.6 Metrics
We evaluate the models using three key metrics across our five datasets.
3.6.1 Bias Dataset Effectiveness Metrics
We measure both performance and fairness for three of the five datasets. For VisoGender and VLStereoSet, we adopt the evaluation metrics provided by those datasets.
For the portrait datasets (UTKFace, CelebA, and PATA), we use the Macro-F1 score to assess performance. To evaluate fairness, we employ the Demographic Parity Ratio (DPR) implemented in fairlearn Bird et al. (2020). DPR measures the ratio between the highest and lowest selection rates. The selection rate is defined as the predicted positive rate for a given group/protected category. In a fair model, this selection rate should be independent of the underlying group/protected category, meaning the DPR should ideally be close to 100%.
For VisoGender, we use the provided metric: Resolution Bias (RB). Resolution Bias is the difference between the number of correctly resolved pronouns for males and females within a specific occupation. A fair model should have a Resolution Bias of 0 for each occupation, indicating equal performance across genders. A positive score means that it is more accurate for females and a negative score means that it is more accurate for males. To compute the overall RB, we average the values across all occupations.
For VLStereoSet, we use the provided Vision-Language Bias Score (VLBS). VLBS is defined as the percentage of instances where the model selects a stereotypical caption over an anti-stereotypical caption. A completely unbiased model would yield a VLBS of 0.
4 Results
Figure 3: Performance of models on all the portrait datasets compared to their fairness (DPR = demographic parity ratio, higher the better.). The dashed diagonal line represents the dividing line between trading off performance for fairness or vice versa. Arrows indicate changes in performance and fairness from CLIP to LLaVa models when the differences for both are > 10%.
4.1 Portrait Datasets
Adding the LLM on top of ViT improves performance/fairness tradeoffs.
The results of our evaluation are presented in Figure 3. Upon closer examination, we observe a significant performance gap between the Vision Language Models (VLMs) and the baseline CLIP models. Across all datasets, the VLMs either match or outperform the CLIP models in terms of Macro-F1 score. For example, in the CelebA dataset, we see a 30-point improvement in predicting heavy makeup and a 2-point increase in predicting blondness. In the UTKFace dataset, we observe a 20-point performance boost for predicting age, a 30-point improvement for predicting race, and a modest 5-point increase for predicting gender.
Fairness does not generalize consistently across tasks.
Although integrating LLMs with CLIP models generally enhances both fairness and performance for most datasets, the tradeoffs between these two factors are inconsistent. For instance, in the CelebA blond hair dataset, we observe improvements in fairness ranging from 2 to 20 points. For UTKFace Age, we see a 20-point fairness gain when measuring performance across races. For UTKFace Race, a 4-6 point improvement across ages. In UTKFace Gender, we also see a 25-point increase in fairness when measuring across races.
However, these improvements are not universal. As mentioned, CelebA heavy makeup and UTKFace Age Gender exhibit a troubling reduction in fairness, raising the concern that fairness for one task does not reliably generalize to other tasks, even within the same image dataset. Practitioners should be cautious when assuming that fairness achieved in one domain will extend to others, even when they use the same image data.
4.2 Scene-Based Bias Evaluation
4.2.1 VLStereoSet
VLStereoSet may be inadequate for studying image-induced bias in VLMs.
As shown in Figure 5 (see Appendix), the results suggest that VLStereoSet struggles to effectively evaluate whether images elicit bias in Vision Language Models (VLMs). Without the image input, all models exhibit a Vision-Language Bias Score (VLBS) around 0.5-0.6, indicating a significant level of bias since lower scores are better, and random chance would yield a score of 0.33. Including the image results in only a marginal improvement in fairness, with a bias reduction of approximately 5 points for three models and 15 points for LLaVa Vicuna 7B. This finding highlights an issue with VLStereoSet as a dataset. The modest performance improvements compared to gains observed in other datasets, such as CelebA and UTKFace, suggest that the model can often infer the correct answer without relying on the image.
Figure 4: Average Demographic Parity Ratio comparison between "Text Only" and "With Image" for PATA (left). Macro-F1 comparison between "Text Only" and "With Image" for PATA (right). The grey bar represents the baseline CLIP Large 336 model, which is the backbone frozen vision encoder for all LLaVa that we have tested here.
4.2.2 PATA
PATA may have some limitations when used as a bias evaluation dataset.
As seen in Figure 4, the PATA dataset also exhibits issues in its design, rendering it insufficient for evaluating bias in models. Across all models, performance either decreases or remains unchanged when images are included. This stagnation or drop in performance suggests that PATA does not require models to leverage the image input as intended.
Moreover, the fairness evaluation (Figure 4) reveals that models tend to become less fair or show no improvement when images are provided. The fact that models can achieve nearly perfect fairness using text-only inputs suggests that the inclusion of images is irrelevant and may even confuse the models. This undermines the purpose of PATA as a scene-based bias evaluation dataset, where the images should play a critical role in model assessment. Instead, PATA appears to rely solely on textual cues.
Model Dataset Acc. ↑ Average RB ↓ CF Acc. ↑
Gemini 1.5 Flash $O$ 49.56 -0.9209 0.00
OP 47.33 -0.9397 21.56
LLaVa 34B $O$ 98.68 -0.0179 80.18
OP 87.76 0.0561 57.14
LLaVa Mistral 7B $O$ 99.12 -0.0026 92.95
OP 81.41 -0.2045 58.50
LLaVa Vicuna 7B $O$ 96.48 -0.0718 62.11
OP 77.10 -0.3431 50.11
Random $O$ 50.00 0.0000 50.00
OP 50.00 0.0000 50.00
Table 2: Results for VisoGender Evaluaton. Acc is accuracy. CF Acc is the accuracy of the adversarial version of the dataset. The performance gap between Acc vs CF Acc shows how robust models are.
4.3 Pronoun Resolution with VisoGender
The original VisoGender dataset is insufficient to stress-test model bias.
In Table 2, we compare the performance of our models on both the Occupation-Object (OO) and Occupation-Participant (OP) versions of the VisoGender dataset. For the OO subset, all models, except Gemini 1.5 Flash, achieve near-perfect accuracy. On the OP dataset, the trend continues, with most models maintaining around 80% accuracy, except for Gemini 1.5 Flash, which performs worse than random. It answers 95% of the time for the female option for both datasets likely due to interference from its safety filtering mechanisms.
Fairness evaluation reveals subtle biases in the OP dataset.
For the OO subset, most models exhibit near-unbiased performance, with bias scores close to zero. However, the OP dataset reveals more nuanced biases. Specifically, the 7B LLaVa models tend to be more accurate for images where females are the primary focus, while LLaVa 34B exhibits an opposite bias, showing better accuracy for images with males.
The OP dataset provides more valuable insights into model bias than the OO subset.
When measuring the models’ ability to reveal biases, the OO dataset proves ineffective, as all LLaVa models achieve near-perfect accuracy and fairness, or, in the case of Gemini, fail to answer adequately. In contrast, the OP dataset provides more informative results: LLaVa Mistral and Vicuna models demonstrate a bias towards females, while LLaVa 34B is slightly biased toward males. These insights suggest that the OP dataset is far more useful for uncovering gender-related biases in pronoun resolution tasks than the OO version.
4.4 Adversarial Dataset for Pronoun Resolution
Adversarial prompts reveal models’ reliance on textual cues over visual information.
As shown in Table 2, model performance drops significantly when adversarial prompts are introduced, indicating a reliance on textual information even when it conflicts with visual cues. Most models experience a 20-30 point drop in accuracy from their baseline performance. Notably, LLaVa Mistral 7B shows the least degradation, with only a 6-point drop in accuracy on the OO dataset, suggesting it is more robust to adversarial prompts than the other models. On the other hand, Gemini 1.5 Flash suffers the most, with a 50-point accuracy drop on the OO dataset and a 25-point drop on the OP dataset, despite being one of the largest and most powerful models tested.
These results highlight the brittleness of models when visual and textual information conflict.
The significant accuracy drops across almost all models suggest that they tend to prioritize textual information over visual cues when the two are in conflict. This phenomenon, often referred to as "sycophancy," exposes a key weakness in current bias evaluation methods: models are highly susceptible to misleading text, which skews the bias analysis. To ensure accurate assessments of bias, datasets and prompts must be meticulously designed to avoid introducing conflicting or extraneous information.
The adversarial dataset serves as a more rigorous benchmark for bias evaluation.
The introduction of adversarial prompts shows that most models are less robust when the visual content is at odds with the textual description, underscoring the importance of well-designed datasets for bias evaluation. The adversarial version of VisoGender provides a much-needed stress test, revealing hidden biases that are not evident in the original dataset.
5 Discussion and Conclusion
This paper presents one of the first comprehensive evaluations of bias in VLMs, analyzing state-of-the-art models across major datasets to assess both model bias and the effectiveness of bias evaluation techniques.
‘Portrait’ datasets are effective for bias evaluation.
Portrait-based datasets like UTKFace and CelebA effectively highlight fairness gaps between LLaVa and CLIP models, showing that VLMs are generally less biased and more performant when paired with an LLM backbone. However, biases in these models vary significantly depending on the dataset, emphasizing the need for diverse datasets to capture different bias dimensions.
‘Scene’ datasets fail to accurately measure bias.
Scene-based datasets like VLStereoSet and PATA suffer from data artifacts that allow models to exploit text cues, underestimating bias. Future work could refine these datasets to challenge models to properly integrate visual and textual data.
‘Limitations of VisoGender’
Bias evaluation in VLMs is particularly challenging due to models often prioritizing textual cues over visual information, a behavior known as sycophancy. This tendency leads to biased outputs when text and images conflict. The original VisoGender dataset is no longer an effective benchmark, as models now achieve near-perfect accuracy, suggesting it does not to truly test a modern model’s ability to integrate visual information. Instead, models rely heavily on text, masking underlying biases. The adversarial version of VisoGender reveals more gender-related biases, emphasizing the potential value and need for more rigorous and carefully designed datasets that challenge models to "see" and resist being misled by textual inputs.
In conclusion, portrait datasets currently offer the best tools for bias detection, while scene-based and pronoun resolution datasets require significant improvements. Future work should focus on developing more rigorous and carefully designed datasets to ensure VLMs are both fair and reliable across a variety of tasks.
6 Limitations
We identify the following limitations in our work. First, due to the growing popularity of multimodal models, it is impractical to evaluate all available models. We have focused on a representative subset acknowledging that this might limit the generalizability of our findings. Second, some of the images for VisoGender, PATA, and VLStereoSet were unable to be retrieved due to broken hyperlinks, possibly affecting the quality of the evaluations. Third, given the limited datasets available, we recognize that the models’ fairness discussed in this paper may not fully not extend to real-world scenarios.
Acknowledgments
We would like to thank UAB Research Computing for use of their HPC infrastructure. The authors acknowledge financial support from the Google PhD Fellowship (SC), the Woods Foundation (DB, SC), the NIH (NIH-USA R01CA294033 (DB, SC), NIH-USA U54CA274516-01A1 (DB, SC), the ASTRO-ACS Clinician Scientist Development Grant ASTRO-CSDG-24-1244514 (DB). The authors also thank Google Cloud research fund for Gemini API inference costs.